OpenStudy (anonymous):

The sum of 237 and six times the opposite of a number is 405. What is the number? A) -28 B) -24 C) 30 D) 32

OpenStudy (mathstudent55):

Let the unknown number be x. Using x, how do you write the opposite of a number?

OpenStudy (anonymous):

a negative number?

OpenStudy (mathstudent55):

Yes, it's the negative of x, or simply -x

OpenStudy (anonymous):

6x

OpenStudy (mathstudent55):

Now we have the opposite of a number is -x. How do you write "six times the opposite of a number"?

OpenStudy (mathstudent55):

Be careful. 6x is simply "six times a number." We need 6 times the opposite of a number.

OpenStudy (anonymous):

6x+n?

OpenStudy (mathstudent55):

The opposite of a number is -x. We need six times -x

OpenStudy (mathstudent55):

We are not adding anything yet. six times a number is 6x the opposite of a number is -x six times the opposite of a number is 6 * (-x) ok so far?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

OK. Now we deal with 237. The "sum" means an addition. The sum of 237 and six times the opposite of a number is 237 + 6 * (-x) Ok?

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

Now we finally take care of the 405. 237 + 6 * (-x) = 405 Now we have an equation. We can solve the equation and find the number.

OpenStudy (mathstudent55):

First, let's simplify 6 * (-x) into simply -6x 237 - 6x = 405 Now we subtract 237 from both sides.

OpenStudy (mathstudent55):

237 - 6x = 405 -237 -237 ---------------- -6x = 168

OpenStudy (mathstudent55):

Now we divide both sides by -6 -6x/(-6) = 168/(-6) x = -28

OpenStudy (mathstudent55):

The number is -28.

OpenStudy (anonymous):

yup that exactly what i got

OpenStudy (mathstudent55):

Great.

OpenStudy (anonymous):

thanks for the help really appreciated that